Scores were obtained from a hypothetical test A. Only 12 scores from a large group in which the distribution of scores is approximately normal are presented here.

75 61 93 84 11 50 36 42 57 76 46 70

Mean=58

SD=23

58.41667 23.00774


(1) What are the Z values for the score 42 and 75?
Answer #1

Answer)

Z = (x - mean)/s.d

Given mean = 58

S.d = 23

For 42

Z = (42 - 58)/23 = -0.7

For 75

Z = (75 - 58)/23 = 0.74